# Drivel-ology: Challenging LLMs with Interpreting Nonsense with Depth

### v1

@@ 3.3 Data Annotation
Labelling Drivelology requires both content comprehension and cultural context understanding. We implemented a rigorous four-step annotation protocol: (1) Annotator Selection.  We assembled a team of seven multilingual annotators [^3] The annotation team consisted of four authors of this paper and three paid annotators recruited for their linguistic expertise. who all held at least a Master’s degree and demonstrated proficiency in multiple languages. (2) Drivelology Detection and Tagging.  Annotators identified texts as either Drivelology or non-Drivelology, and classified Drivelology samples into categories including Misdirection, Paradox, Switchbait, Inversion, and Wordplay. (3) Implicit Narrative Writing.  We employed a human-in-the-loop process to create the narrative descriptions. For each Drivelology sample, human experts drafted and refined the correct narrative explanation. We then utilised GPT-4.5444gpt-4.5-preview-2025-02-27 as an assistive tool to generate four plausible but incorrect narrative counterparts, all of which underwent a final stage of manual verification and editing to ensure their quality as effective distractors. (4) Quality Check.  A meta-reviewer with linguistics and psychology expertise reviewed all annotations, excluding ambiguous or controversial samples. The meta-reviewer also revised the narratives as needed to ensure consistent length, uniform writing style, and improve overall readability. Further details about the annotation process can be found in Appendix A.1.
@@ 3.4 Task Design
Drivelology Detection.   A binary classification task where the model must determine if a given text is “Drivelology” or “non-Drivelology.”
Drivelology Tagging.   A multi-label classification task. The model assigns one or more descriptive categories (see §3.1) to a Drivelology sample to capture its layered rhetorical structure.
Narrative Writing.   A generative task that assesses the model’s ability to articulate the coherent implicit narrative and underlying meaning of a Drivelology sample, requiring it to move beyond a surface-level reading.
Narrative Selection.   This multiple-choice question answering (MCQA) task asks the model to pick the correct narrative for a Drivelology sample from several options. The Easy version offers one correct answer and four distractors, which we generated by prompting GPT-4.5 to produce plausible but incorrect interpretations of the sample. These AI-generated options were then manually reviewed and refined by human experts to ensure they were effective and challenging. The Hard version adds a “none of the above” option, requiring deeper reasoning, as this option should only be chosen if none of the provided narratives adequately capture the underlying meaning of the Drivelology sample. This additional step significantly increases the task’s complexity, as it prevents reliance on simple elimination strategies.
@@ 4.1 Models and Settings
We evaluate the models’ performance in a zero-shot setting using recent LLMs. We include both proprietary models such as GPT-4 Achiam et al. (2023) and Claude-3 Anthropic (2024), as well as open-sourced models including Qwen3 Qwen Team (2025), Qwen2.5 Qwen Team (2024), Llama3.1 Grattafiori et al. (2024), Llama3 Grattafiori et al. (2024), and DeepSeek V3 Liu et al. (2024a).
@@ 4.2 Evaluation Metrics
For the understanding tasks, we use accuracy for Drivelology Detection task, F1 score for Drivelology Tagging task, and accuracy for the MCQA task. For the generation task that involves narrative writing, we apply reference-based evaluation metrics commonly used in text generation studies Celikyilmaz et al. (2020). Specifically, we use BERTScore Zhang et al. (2020) and LLM-as-a-judge Zheng et al. (2023). Recent work shows that GPT-based evaluation aligns well with human judgments Chan et al. (2023); Liu et al. (2023); Hu et al. (2024); Gu et al. (2024), thus we apply GPT-4 for LLM-as-a-judge evaluation. Note that we use different GPT variants for different purposes: gpt-4.5 for data annotation, gpt-4o-mini for zero-shot experiments, and gpt-4.1 for LLM-as-a-judge evaluation in text generation tasks. This helps reduce potential evaluation bias toward GPT-4’s own generation.
@@ 5 Main Results
Table 1: Main results. For narrative writing task, we report BERTScore-recall (BERT) and GPT-4-as-a-judge (GPT) evaluation scores. For narrative selection task, we report accuracy. For Drivelology classification tasks, we report accuracy for detection task and report weighted F1 score for tagging task. Best scores are bold and the second best ones are marked with underline.
The main results (Table 1) show a clear hierarchy in model performance. Deepseek-v3 is the dominant model, achieving the top score in five of the six evaluated metrics. The contrast between the two evaluation metrics in the Narrative Writing task is particularly noteworthy. While BERTScore-recall values are high across all models (ranging from 84.67% to 87.11%), suggesting a universal proficiency in generating fluent text, the GPT-4-as-a-judge scores (rated on a 1-5 Likert scale) provide a much clearer picture of true narrative quality. On this scale, deepseek-v3 (3.59) and claude-3.5-haiku (3.39) are the only models to score comfortably above three, indicating their outputs were judged as possessing high semantic quality. In stark contrast, other models like llama-3-8b-instruct (2.63) and qwen3-8b-instruct (2.64) fall below this threshold, suggesting their narratives failed to capture the required depth and were deemed qualitatively weaker by the LLM-as-a-judge. The most striking performance gap presents in the MCQA task. The Hard setting causes a steep decline in accuracy for all models, exposing a critical weakness in subtle reasoning. Notably, qwen3-8b-instruct is a notable outlier here, scoring 26.78%, which far surpasses the next-best model. In the Classification tasks, deepseek-v3 again confirms its superior understanding by leading in both Detection (81.67%) and Tagging (55.32%) tasks.
@@ 5.1 Prompt Language Influence
An analysis of the impact of prompt language on model performance reveals a metric-dependent pattern. As shown in Figure 2, the choice between English and Mandarin prompts is not a neutral design but a significant factor that influences evaluation outcomes. We identify two distinct and opposing patterns. First, English prompts consistently yield superior performance on tasks that reward lexical precision and complex logical reasoning. In the Narrative Writing task, this is most evident in the BERTScore results, where every model scores higher when prompted in English. This suggests that English instructions may prime the models to generate outputs with greater lexical overlap with the reference translations, a feature to which BERTScore is highly sensitive Hanna and Bojar (2021). A similar advantage for English prompts is observed in MCQA task. The uniform improvement in both Easy and Hard settings implies that English may serve as a more robust internal language of reasoning. Conversely, Mandarin prompts produce a consistent, albeit smaller, advantage on tasks that prioritise direct content comprehension. This pattern holds across three metrics: the GPT-as-a-judge scores for Narrative Writing, and both Detection and Tagging tasks. The improved performance under GPT-as-a-judge, which evaluates qualitative coherence, indicates that Mandarin prompts better align the models with the semantic and narrative intent of the source material. The consistent gains in the classification tasks further suggest that for direct comprehension and categorisation, instructions in Mandarin are more effective.
5.2 Role of Language in MCQA Task
@@
A closer look at the Narrative Selection (MCQA) results from Table 1 reveals that aggregate scores mask significant performance variations across the different languages in our DrivelHub dataset. As shown in the breakdown in Figure 3, we can analyse the difficulty of each language’s content for the models. Among the models, deepseek-v3 consistently demonstrates the most robust cross-lingual performance, achieving the highest accuracy across nearly all languages in both the Easy and Hard settings. The analysis also pinpoints which languages pose a greater challenge. Korean and Mandarin consistently result in the lowest accuracy, especially in the Hard task, marking their content as the most difficult for the models to process.
5.3 Model Size Scaling on Narrative Selection
Prompt Qwen3 Size MCQA
Easy Hard
English 4B 81.00 6.00
8B 83.17 26.78
14B 83.94 45.83
Mandarin 4B 77.61 2.44
8B 81.11 19.11
14B 83.50 47.89
Table 2: MCQA performance of Qwen3 models. This table shows the accuracy on the Easy and Hard tasks when prompted in English and Mandarin. Full version of all tasks can be checked in Table 5.
The results in Table 2 clearly demonstrate a positive scaling trend for the Qwen3 series, particularly on the more challenging Hard version of the MCQA task. For the Easy task, performance gains are modest but consistent: as model size increases from 4B to 14B, accuracy improves by approximately 3% for English prompts and 6% for Mandarin prompts. However, the Hard task reveals a much more spiking trend. Here, performance scales exceptionally well with model size. With an English prompt, the score leaps from a mere 6.00% for the 4B model to 45.83% for the 14B model. The effect is even more pronounced with a Mandarin prompt, where the score skyrockets from 2.44% to 47.89%. This indicates that the more complex reasoning required by the Hard task is a key differentiator that is unlocked by larger model sizes. Therefore, the ability to handle such complex reasoning appears to be an emergent property in the Qwen3 architecture, strongly correlated with its parameter count.
@@ 6 Analysis and Discussion
6.1 Analysis in Model Reasoning
In Narrative Writing task, claude-3.5-haiku and deepseek-v3 achieve the highest GPT-4-as-a-judge scores (3.39 and 3.59) and also perform strongly in Drivelology Tagging task (52.03% and 55.32%). This correlation between their performance in both tasks raises an important question: Do these models arrive at correct Drivelology classifications through appropriate reasoning that reflects a true understanding of the underlying meaning? To investigate this question, we analyse their reasoning processes across several representative examples.
For example: “Meng Po: Those who have forgotten their names, please follow me.” Deepseek-v3 categorise this as switchbait, emphasising the cultural significance of Meng Po, a mythological figure who administers the Soup of Forgetfulness in Chinese folklore. Its reasoning explicitly highlights the importance of cultural knowledge, suggesting they treat Meng Po’s mythological role as important context that readers must understand to appreciate the Drivelology. In contrast, claude-3.5-haiku categorises it as an paradox, focusing on the logically self-contradictory statement: “how can someone who has forgotten their name respond to such a call?” This divergence in reasoning approaches suggests varying degrees of cultural knowledge internalisation among models. Claude-3.5-haiku appears to have so thoroughly internalised the cultural context of Meng Po that it treats it as implicit knowledge, allowing it to focus on the logical structure of the text rather than its cultural elements. This observation raises important questions about how different models process and prioritise cultural knowledge versus logical reasoning in their analysis of culturally-embedded texts, and how such internalisation affects their ability to identify different categories of Drivelology.
6.2 Analysis in Human Reasoning
@@
Consider the statement: “I hate two kinds of people: the first kind is those who don’t finish their…”. From a paradox perspective, the speaker claims to dislike people who speak incompletely, yet the sentence itself is left unfinished, ironically exemplifying the very behaviour it criticises. This self-contradiction highlights the speaker’s insincerity and creates a paradoxical effect. Alternatively, from a misdirection viewpoint, the statement sets up the expectation of a complete list, but then abruptly stops, leaving the audience anticipating an answer that never comes. The humour and irony arise from this unresolved expectation. Another example is: “I deeply admire Che Guevara’s anti-capitalist spirit, so I bought all his merchandise.” Here, the paradox lies in admiring Guevara’s anti-capitalist stance while simultaneously engaging in capitalist consumerism by buying his merchandise. This contradiction turns ideological admiration into commercial participation. The switchbait interpretation depends on cultural knowledge: recognising Che Guevara as a symbol of anti-capitalism is key. Without this context, the contradiction, and the humour, may not be apparent. The text’s irony and layered meaning rely on shared cultural and historical understanding, making switchbait also an appropriate label.
@@ 7 Conclusion
In this work, we introduce Drivelology, a unique linguistic phenomenon that challenges the semantic and pragmatic understanding of LLMs. By constructing and evaluating the DrivelHub dataset across multiple languages and task settings, we demonstrate that current LLMs, despite their strong ability, still struggle to interpret and reason about texts that blend surface-level nonsense with deeper, implicit meaning. Our findings highlight the need for more advanced models capable of non-linear comprehension beyond statistical pattern-matching. We hope that our dataset and benchmarks will inspire further research into the modelling of creative and contextually rich language, ultimately advancing the development of AI systems with deeper social and cultural awareness.
@@
Language Imbalanced.  Almost half of the samples in the DrivelHub dataset are in Mandarin, while the remaining samples are distributed among English, Spanish, French, Japanese, and Korean. This results in a slight language imbalance, which may affect the generalisability of our findings across different linguistic and cultural contexts. We do our best in §5 to control for potential content distribution bias arising from this imbalance. Additionally, as the DrivelHub dataset is still being expanded, we will continue to focus on addressing distribution differences by increasing the representation of Drivelology samples in underrepresented languages.
Limited Computation Resources.  Due to budget constraints, we were unable to evaluate stronger proprietary LLMs such as GPT-4.5, Claude-3.7, or DeepSeek R1, as their usage costs are prohibitively high. For open-source models, we restricted our experiments to 14B parameter models because of limited computational resources, and were unable to run larger models within our available infrastructure. We encourage researchers and the broader community to expand on this work by evaluating larger or more advanced LLMs as resources permit.
Focus on Understanding Rather Than Generation.  In this paper, we focus on evaluating the understanding and reasoning abilities of LLMs with respect to Drivelology, rather than their capacity to generate fluent and human-like Drivelology text. While generation is an important aspect, it falls outside the main scope of our study. Nevertheless, we include a discussion in the Appendix C with sample generations, illustrating that current LLMs often require over 20 attempts to produce a Drivelology text that achieves comprehensive alignment between topic, rhetorical category, and sentence structure.

### v2

@@ 3.3 Data Annotation
Labelling Drivelology requires both content comprehension and cultural context understanding. We implemented a rigorous four-step annotation protocol: (1) Annotator Selection:  We assembled a team of seven multilingual annotators who all held at least a Master’s degree and demonstrated proficiency in multiple languages. [^4] The annotation team consisted of four authors of this paper and three paid annotators recruited for their linguistic expertise. (2) Drivelology Detection and Tagging:  Annotators identified texts as either Drivelology or non-Drivelology, and classified Drivelology samples into categories including Misdirection, Paradox, Switchbait, Inversion, and Wordplay. (3) Implicit Narrative Writing:  We employed a human-in-the-loop process to create the narrative explanations. For each Drivelology sample, human experts drafted and refined the correct narrative explanation. We then utilised GPT-4.5555gpt-4.5-preview-2025-02-27 as an assistive tool to generate four plausible but incorrect narrative counterparts, all of which underwent a final stage of manual verification and editing to ensure their quality as effective distractors. (4) Quality Check:  A meta-reviewer with linguistics and psychology expertise reviewed all annotations. The meta-reviewer also revised the narratives as needed to ensure consistent length, uniform writing style, and improve overall readability. Further details concerning the annotation process can be found in Appendix A.1.
@@ 3.4 Task Design
Drivelology Detection.   A binary classification task where the model must determine if a given text is Drivelology or non-Drivelology.
Drivelology Tagging.   A multi-label classification task where the model assigns one or more descriptive categories (see §3.1) to a Drivelology sample to capture its layered rhetorical structure.
Narrative Writing.   A generative task where the model explains the implicit narrative and underlying meaning of a Drivelology sample, requiring it to move beyond a surface-level reading.
Narrative Selection.   A multiple-choice question answering (MCQA) task where the model is tasked with selecting the correct narrative for a Drivelology sample from several options. The Easy version offers one correct answer and four distractor, whilst the Hard version adds a “none of the above” option, requiring deeper reasoning, as this option should only be chosen if none of the provided narratives adequately capture the underlying meaning of the Drivelology sample. This additional step significantly increases the task’s complexity, as it prevents reliance on simple elimination strategies.
@@ 4.1 Models and Settings
We evaluate the performance of state-of-the-art LLMs in a zero-shot setting. We utilise both proprietary models including GPT-4 Achiam et al. (2023) and Claude-3 Anthropic (2024), as well as open-sourced models including Qwen3 Qwen Team (2025), Qwen2.5 Qwen Team (2024), Llama3.1 Grattafiori et al. (2024), Llama3 Grattafiori et al. (2024), and DeepSeek V3 Liu et al. (2024a).
@@ 4.2 Evaluation Metrics
We use accuracy for the Drivelology Detection task, F1 score for the Drivelology Tagging task, and accuracy for the MCQA task. For the generation task that involves writing narrative explanations, we apply reference-based evaluation metrics commonly used in text generation studies Celikyilmaz et al. (2020). Specifically, we use BERTScore Zhang et al. (2020) and an LLM-as-a-judge evaluation paradigm Zheng et al. (2023). Recent work shows that GPT-based evaluation aligns well with human judgments Chan et al. (2023); Liu et al. (2023); Hu et al. (2024); Gu et al. (2024), and thus we select GPT-4 series for LLM-as-a-judge evaluation. The judge was tasked to rate each generated narrative on a 1-to-5 Likert scale based on its semantic quality. Note that we use different GPT variants for different purposes: gpt-4.5 for data annotation, gpt-4o-mini for zero-shot experiments, and gpt-4.1 for LLM-as-a-judge evaluation in text generation tasks. This helps reduce potential evaluation bias toward GPT-4’s own generation Hu et al. (2024); Liu et al. (2024b).
@@ 5 Main Results
deepseek-v3 87.11 3.59 86.83 15.50 81.67 55.32
Table 1: Main results. For the narrative explanation writing task, we report BERTScore-recall (BERT) and GPT-4-as-a-judge (GPT) evaluation scores. For the narrative selection task, we report accuracy. For the Drivelology classification tasks, we report accuracy for the detection task and weighted F1 score for the tagging task. The best scores are bold and the second best ones are underlined.
The main results in Table 1 show a clear hierarchy in model performance. Deepseek-v3 is the dominant model, achieving the top score in five of the six evaluated metrics. The contrast between the two evaluation metrics in the Narrative Writing task is particularly noteworthy. While BERTScore-recall values are high across all models, suggesting a universal proficiency in generating fluent text, the GPT-4-as-a-judge scores provide a much clearer picture of true narrative quality. On this scale, deepseek-v3 (3.59) and claude-3.5-haiku (3.39) are the only models to score comfortably above three, indicating their outputs were judged as possessing high semantic quality. In stark contrast, other models like llama-3-8b-instruct (2.63) and qwen3-8b-instruct (2.64) fall below this threshold, suggesting their narratives failed to capture the required depth and were deemed qualitatively weaker by the LLM-as-a-judge. The most striking performance gap is present in the MCQA task. The Hard setting causes a steep decline in accuracy for all models, exposing a critical weakness in subtle reasoning. Notably, qwen3-8b-instruct is an outlier here, scoring 26.78%, which far surpasses the next-best model. In the Classification tasks, deepseek-v3 again confirms its superior understanding by leading in both the Detection (81.67%) and Tagging (55.32%) tasks.
@@ 5.1 Prompt Language Influence
An analysis of the impact of prompt language on model performance reveals a metric-dependent pattern. As shown in Figure 3, the choice between English and Mandarin prompts is not a neutral choice but a significant factor that influences evaluation outcomes. We identify two distinct and opposing patterns. Firstly, English prompts consistently yield superior performance on tasks that reward lexical precision and complex logical reasoning. In the Narrative Writing task, this is most evident in the BERTScore results, where every model scores higher when prompted in English. This suggests that English instructions may prime the models to generate outputs with greater lexical overlap with the reference translations, a feature to which BERTScore is highly sensitive Hanna and Bojar (2021). A similar advantage for English prompts is observed in the MCQA task. The uniform improvement in both Easy and Hard settings implies that English may serve as a more robust internal language of reasoning. Conversely, Mandarin prompts produce a consistent, albeit smaller, advantage on tasks that prioritise direct content comprehension. The improved performance under GPT-as-a-judge, which evaluates qualitative coherence, indicates that Mandarin prompts better align the models with the semantic and narrative intent of the source material. The consistent gains in the classification tasks further suggest that for direct comprehension and categorisation, instructions in Mandarin are more effective.
5.2 Model Size Scaling in the Qwen3 Series
The results in Table 2 illustrate the impact of model size on performance. We focus on the Qwen3 series across the MCQA and Classification tasks, as these tasks exhibited the widest performance variance in Table 1, making them most suitable for studying scaling effects. For the Easy MCQA task, performance gains are consistent but modest: as model size increases from 4B to 14B, accuracy improves by approximately 3% for English prompts and 6% for Mandarin. The Hard task reveals a spiking scaling effect. With English prompts, accuracy leaps from a mere 6.00% for the 4B model to 45.83% for the 14B model. The trend is even more pronounced with Mandarin prompts, where the score skyrockets from 2.44% to 47.89%. This indicates that the more complex reasoning required by the Hard task is a key differentiator that is unlocked by larger model sizes. Therefore, the ability to handle such complex reasoning appears to be an emergent property in the Qwen3 architecture, strongly correlated with its parameter count. For the Detection task, performance does not consistently improve with size. Notably, when prompted in Mandarin, the 8B model significantly outperforms both its smaller and larger counterparts. The Tagging task reveals yet another pattern: a noticeable dip in performance at the 8B size, which then recovers to achieve the best score at 14B for both languages. These findings indicate that the benefits of model scaling are task-dependent.
Prompt Size MCQA Classification
Easy Hard Detect Tag
English 4B 81.00 6.00 66.80 43.21
8B 83.17 26.78 65.00 38.04
14B 83.94 45.83 66.22 47.61
Mandarin 4B 77.61 2.44 62.86 46.10
8B 81.11 19.11 78.81 41.71
14B 83.50 47.89 71.78 49.13
Table 2: MCQA and Classification results in the Qwen3 series of different sizes. This table shows the performance on both tasks when prompted in English and Mandarin. The Full version containing all tasks can be found in Table 5.
5.3 Role of Language in the MCQA Task
@@
A closer look at the MCQA results from Table 1 reveals that aggregate scores mask significant performance variations across the different languages in DrivelHub. As shown in the breakdown in Figure 4, we can analyse the difficulty of each language’s content for the models. Deepseek-v3 consistently demonstrates the most robust cross-lingual performance, achieving the highest accuracy across nearly all languages in both the Easy and Hard settings. The analysis also pinpoints which languages pose a greater challenge. Korean and Mandarin consistently result in the lowest accuracy, especially in the Hard task, marking their content as the most difficult for the models to process.
@@ 6 Analysis and Discussion
6.1 Analysis of Model Reasoning
In the Narrative Writing task, claude-3.5-haiku and deepseek-v3 achieve the highest GPT-4-as-a-judge scores (3.39 and 3.59) and also perform strongly in the Drivelology Tagging task (52.03% and 55.32%). Thie correlation between their performance in both tasks raises an important question: Do these models arrive at correct Drivelology classifications through appropriate reasoning that reflects a true understanding of the underlying meaning? To investigate this question, we analyse their reasoning processes across several representative examples.
For example: “Meng Po: Those who have forgotten their names, please follow me.” Deepseek-v3 categorise this as switchbait, emphasising the cultural significance of Meng Po, a mythological figure who administers the Soup of Forgetfulness in Chinese folklore. Its reasoning explicitly highlights the importance of cultural knowledge, suggesting they treat Meng Po’s mythological role as important context that readers must understand to appreciate the Drivelology. In contrast, claude-3.5-haiku categorises it as a paradox, focusing on the logically self-contradictory statement: “how can someone who has forgotten their name respond to such a call?” This divergence in reasoning approaches suggests varying degrees of cultural knowledge internalisation among models. Claude-3.5-haiku appears to have so thoroughly internalised the cultural context of Meng Po that it treats it as implicit knowledge, allowing it to focus on the logical structure of the text rather than its cultural elements. This observation raises important questions about how different models process and prioritise cultural knowledge versus logical reasoning in their analysis of culturally-embedded texts, and how such internalisation affects their ability to identify different categories of Drivelology.
6.2 Analysis of Human Reasoning
@@
Consider the statement: “I hate two kinds of people: the first kind is those who don’t finish their…” From a paradox perspective, the speaker claims to dislike people who speak incompletely, yet the sentence itself is left unfinished, ironically exemplifying the very behaviour it criticises. This self-contradiction highlights the speaker’s insincerity and creates a paradoxical effect. Alternatively, from a misdirection viewpoint, the statement sets up the expectation of a complete list, but then abruptly stops, leaving the audience anticipating an answer that never comes. The humour and irony arise from this unresolved expectation. Another example is: “I deeply admire Che Guevara’s anti-capitalist spirit, so I bought all his merchandise.” Here, the paradox lies in admiring Guevara’s anti-capitalist stance while simultaneously engaging in capitalist consumerism by buying his merchandise. This contradiction turns ideological admiration into commercial participation. The switchbait interpretation depends on cultural knowledge: recognising Che Guevara as a symbol of anti-capitalism is key. Without this context, the contradiction, and the humour, may not be apparent. The text’s irony and layered meaning rely on shared cultural and historical understanding, making switchbait also an appropriate label.
@@ 7 Conclusion
In this work, we introduced Drivelology, a unique linguistic phenomenon that challenges the semantic and pragmatic understanding of LLMs. We constructed and evaluated the DrivelHub dataset across multiple languages and task settings. Our extensive experiments reveal a critical and consistent gap between statistical fluency and genuine comprehension in state-of-the-art LLMs. While models can generate syntactically coherent text, they largely fail to grasp the layered, culturally-embedded meanings central to Drivelology. We found that complex reasoning, particularly on the “Hard” MCQA task, remains a significant bottleneck, though performance scales predictably with model size. Conversely, performance on classification tasks showed non-linear scaling, suggesting that simply increasing parameter count is not a panacea for all reasoning deficits. The failure of LLMs to interpret Drivelology underscores a deep representational gap in their ability to model complex social and cultural contexts. Our work provides a concrete benchmark for the community to address these deeper challenges. Future research should focus not only on scaling models but also on developing novel training paradigms that explicitly target the multi-layered reasoning that defines sophisticated human communication.
@@
Language Imbalance.  Over one third of the samples in the DrivelHub dataset are in Mandarin (see Table 4). This results in a slight language imbalance, which may affect the generalisability of our findings across different linguistic and cultural contexts. We do our best in §5 to control for potential content distribution bias arising from this imbalance. Additionally, as the DrivelHub dataset is still being expanded, we will continue to focus on addressing distribution differences by increasing the representation of Drivelology samples in underrepresented languages.
Limited Computation Resources.  Due to budget constraints, we were unable to evaluate stronger proprietary LLMs such as GPT-5, Claude-3.7, or DeepSeek R1, as their usage costs are prohibitively high. For open-source models, we restricted our experiments to 14B parameter models because of limited computational resources, and were unable to run larger models within our available infrastructure. We encourage researchers and the broader community to expand on this work by evaluating larger or more advanced LLMs as resources permit.
Focus on Understanding Rather Than Generation.  In this paper, we focus on evaluating the understanding and reasoning abilities of LLMs with respect to Drivelology, rather than their capacity to generate fluent and human-like Drivelology text. While generation is an important aspect, it falls outside the main scope of our study. Nevertheless, we include a discussion in Appendix C with sample generations, illustrating that current LLMs often require over 20 attempts to produce Drivelology that achieves comprehensive alignment between topic, rhetorical category, and sentence structure.